OpenStudy (anonymous):

Write the equation in standard form y=1/3x-2/5

jigglypuff314 (jigglypuff314):

standard form: ax + by = c so you want x and y values on one side by themselves so subtract both sides by 1/3x so -1/3x + y = -2/5 but standard form for a line wants integer coefficients so multiply through by a common multiple to cancel out all denominators so 5x - 15y = 6